OpenStudy (anonymous):

Nico is saving money for his college education. He invests some money at 6%, and 1200 less than that amount at 5%. The investments produced a total of 160 interest in a year. How much did he invest at each rate?

OpenStudy (anonymous):

@cwrw238 Can you please help me with this question

OpenStudy (anonymous):

Let \(x\) be the amount he invested for the first one. The second one he invested \(x-1200\).

OpenStudy (anonymous):

\[ x\times(100\%+6\%) + (x-1200)\times(100\%+5\%) = 160 \]

OpenStudy (anonymous):

You have to simplify this and solve for \(x\).